# Revisiting Randomization in Greedy Model Search

### v1

@@ 3.1 Analytical Framework
The features are uncorrelated with respect to the sample, i.e., $\mathbf{X}^{\top}\mathbf{X}/n=\mathbf{I}$.
While unlikely to describe most real datasets, Assumption 3.1 allows us to derive exact asymptotic expressions for coefficient behavior and precisely characterize how the randomization from feature subsampling propagates through the FS algorithm. This, in turn, enables direct comparisons with the broader regularization landscape, including ridge, lasso [50], elastic net [62], LARS [23], SCAD [25], MCP [60], and SureShrink [20], all of which have been studied in the orthogonal setting.
Remark.
Note that FS, MP, OMP, and BSS are equivalent to each other under Assumption 3.1. In particular, this means that the error reduction that we demonstrate in this section has nothing to do with overcoming the failures of greedy optimization and everything to do with the suboptimal representation obtained via this class of $\ell_{0}$-penalized methods.
3.2 Coefficients and Weights
Let $\hat{f}_{\operatorname{OLS}}=\sum_{j=1}^{p}\hat{\beta}_{j,p}x_{j,p}$ denote the ordinary least squares (OLS) estimator for $f$. Under Assumption 3.1, it is easy to see that its coefficients can be expressed as inner products $\hat{\beta}_{j,p}=\langle\mathbf{y},\mathbf{x}_{\cdot j,p}\rangle$. It turns out that both the FS and RGS estimators can be written in terms of the OLS coefficients. For ease of notation, we establish an ordering convention for the features $x_{1,p},x_{2,p},\ldots,x_{p,p}$ in terms of these coefficients.
@@ 3.2 Coefficients and Weights
The inner products $\hat{\beta}_{j,p}\coloneqq\langle\mathbf{y},\mathbf{x}_{\cdot j,p}\rangle$ are indexed and sorted by magnitude in descending order; that is,
@@
For the sake of simplicity, we assume henceforth that there are no ties in the feature rankings, i.e., the inequalities in (4) are strict. Note also that the definition implies that the feature ordering is random as it depends on the response noise.
Using Definition 3.1 together with the equivalence of FS and BSS, we can write the $k$-step FS estimator as
@@
which recounts the familiar fact that, in the orthogonal setting, FS (and BSS) will select the top-$k$ features most correlated with the response. The indicators $\mathbf{1}(j\leq k)$ can be thought of as weights that modulate the OLS solution. Similarly, the $k$-step RGS estimator has the form
$\hat{f}_{k,m,p}=\sum_{j=1}^{p}w_{j}^{k,m,p}\hat{\beta}_{j,p}x_{j,p}.$ (6)
The RGS weights $w_{1}^{k,m,p},w_{2}^{k,m,p},\ldots,w_{p}^{k,m,p}$ form a non-increasing sequence that takes values between 00 and $1$ and which sum to $k$. Furthermore, they are determined solely by structural parameters ($j,k,m,p$) rather than the actual coefficient values $\hat{\beta}_{j,p}$. In fact, as $m,p\rightarrow\infty$ with a fixed scaling ratio, they are closely approximated by a logistic function with midpoint $k+1/2$ and growth rate $-\log(1-m/p)$, viz.,
@@
This indicates that, in contrast to FS, RGS distributes its weights across a larger number of features through controlled randomization.
We state these facts more rigorously and discuss them further in Section 4. For now, we note that since FS is equivalent to RGS if we set $m=p$, varying $k$ and $m$ in (7) carves out a more flexible family of solutions compared to FS. We will show that fixing $k$ and optimizing over $m$ has the potential to yield both a better fit (via training error) as well as better regularization (via degrees of freedom).
3.3 Degrees of Freedom Improvement
We first show that under a reasonable assumption, decreasing the feature subsampling parameter $m$ (i.e. decreasing the subsampling ratio) has the effect of decreasing the degrees of freedom of the model.
Theorem 3.1.
Grant Assumption 3.1 and suppose $k\mapsto\operatorname{df}(\hat{f}_{k,p})$ forms a concave sequence. [^5] A sequence $c_{1},c_{2},\ldots,c_{p}$ is concave if its first difference sequence $c_{2}-c_{1},c_{3}-c_{2},\ldots,c_{p}-c_{p-1}$ is decreasing. Let $m<m^{\prime}$. Then, for all $k,p\geq 1$, we have
$\operatorname{df}(\hat{f}_{k,m,p})\leq\operatorname{df}(\hat{f}_{k,m^{\prime},p}).$ (8)
In particular, $k\leq\operatorname{df}(\hat{f}_{k,m,p})\leq\operatorname{df}(\hat{f}_{k,p})$ for all $m$.
This result stems from the fact that under Assumption 3.1, the RGS estimator $\hat{f}_{k,m,p}$ can be written as a convex combination of the FS estimators for different values of $k$, while the degrees of freedom operator is linear in its input. The core technical tool is a coupling-based construction of the RGS process, in which estimators at different subsampling levels $m$ are defined on a shared probability space using the same underlying randomness. This allows us to directly compare their induced weight vectors $w_{j}^{k,m,p}$ and establish that they form a majorization hierarchy indexed by $m$ [^6] Formally, for $m<m^{\prime}$, $\sum_{j=1}^{l}w_{j}^{k,m,p}\leq\sum_{j=1}^{l}w_{j}^{k,m^{\prime},p}\quad\text{for all }l=1,\dots,p,$ (9) with equality when $l=p$. This linearity also allows us to argue that the concavity condition in Theorem 3.1 is likely to hold in low SNR scenarios, in particular, when $|\hat{\beta}_{k,p}|\gg|\beta_{k,p}|$. Indeed, in such a scenario, we have
$\begin{split}\frac{\sigma^{2}}{n}\left(\operatorname{df}(\hat{f}_{k,p})-\operatorname{df}(\hat{f}_{k-1,p})\right)&=\frac{\sigma^{2}}{n}\operatorname{df}\big{(}\hat{\beta}_{k,p}x_{k,p}\big{)}\\ &=\mathbb{E}\big{[}\hat{\beta}_{k,p}\big{(}\hat{\beta}_{k,p}-\beta_{k,p}\big{)}\big{]}\\ &\approx\mathbb{E}\big{[}\hat{\beta}_{k,p}^{2}\big{]},\end{split}$ (10)
which is a non-increasing sequence in $k$ according to Definition 3.1. Empirical support for the condition, even in non-orthogonal settings, can be found in Figures 1 and 4 of [32] and Figure 7 of [42].
Remark.
For a constant model (i.e., $\boldsymbol{\beta}=\mathbf{0}$), the approximation in (10) becomes exact.
3.4 Expected Training Error Improvement
Because of the previous section, one can roughly interpret the act of varying $m$ as reallocating the weight/parameter budget of the RGS model among the different features. We now state a lower bound for the reduction in training error attainable from this reallocation. In what follows, we let $\hat{m}$ be the value of $m$ that minimizes the training error $\|\mathbf{y}-\hat{\mathbf{f}}_{k,m,p}\|^{2}$.
Theorem 3.2 (Training Error Improvement).
Grant Assumption 3.1 and fix $k\geq 1$. Assume $p$ is large enough so that $p\geq C\Bigg{(}\frac{\sum_{j=1}^{k}\mathbb{E}\big{[}\hat{\beta}^{2}_{j,p}\big{]}}{\sum_{j=k+1}^{2k}\mathbb{E}\big{[}\hat{\beta}^{2}_{j,p}\big{]}}\Bigg{)}^{3},$ where $C$ is a positive constant that depends only on $k$. Then,
$\mathbb{E}\big{[}\|\mathbf{y}-\hat{\mathbf{f}}_{k,\hat{m},p}\|^{2}\big{]}\leq\mathbb{E}\big{[}\|\mathbf{y}-\hat{\mathbf{f}}_{k,p}\|^{2}\big{]}-\frac{1}{4}\frac{\big{(}\frac{1}{k}\sum_{j=k+1}^{2k}\mathbb{E}\big{[}\hat{\beta}^{2}_{j,p}\big{]}\big{)}^{2}}{\frac{1}{k}\sum_{j=1}^{k}\mathbb{E}\big{[}\hat{\beta}^{2}_{j,p}\big{]}}.$ (11)
Note that the performance gap in (11) is larger when the coefficients $\hat{\beta}_{k+1,p},\hat{\beta}_{k+2,p},\ldots$ corresponding to features omitted by FS has a slower rate of decay. This is unsurprising, because in such a setting, allocating the weight budget more evenly across more features can lead to a better fit.
4 Properties of RGS Weights and Coefficients
In this section, we further analyze the RGS weights $w_{j}^{k,m,p}$ under Assumption 3.1. We first note that the weights are non-increasing. That is, for all $j<j^{\prime}$ and $k,m,p\geq 1$, one can show that $w_{j}^{k,m,p}\geq w_{j^{\prime}}^{k,m,p}$. In addition, we have the following recurrence relation, which will be our main technical tool.
@@ 4 Properties of RGS Weights and Coefficients
$w_{j}^{k,m,p}=\frac{\binom{p-j}{m-1}}{\binom{p}{m}}+\frac{\binom{p-j}{m}}{\binom{p}{m}}w_{j}^{k-1,m,p-1}+$ (12)
$\qquad\qquad\qquad\qquad\qquad\qquad\left(1-\frac{\binom{p-j+1}{m}}{\binom{p}{m}}\right)w_{j-1}^{k-1,m,p-1},$
@@ 4.1 Logistic Approximation
The recurrence (12) is three-dimensional (in $j$, $k$, and $p$) and non-homogeneous, making it highly challenging to find a simple closed-form solution. However, in the scaling limit as $m\to\infty$, $p\to\infty$, and $m/p\to\gamma$ for some constant $\gamma\in[0,1]$, the recurrence converges to a two-dimensional system that decouples into a one-dimensional subsystem with an explicit analytical solution. [^7] Note that it is standard practice to take $m=\lfloor p/3\rfloor$ in RFs, which corresponds to $\gamma=1/3$. This yields the following theorem.
@@
The limiting weights $\tilde{w}_{j}^{k,\gamma}\coloneqq\lim\limits_{\begin{subarray}{c}m\to\infty,\,p\to\infty,\\ m/p\to\gamma\end{subarray}}w_{j}^{k,m,p}$
@@
    where $h(\alpha,k)\coloneqq\frac{1}{\alpha}\log\big{(}e^{\alpha k}-1\big{)}$ and $\alpha\in(0,\infty)$ is defined via $\gamma=1-e^{-\alpha}$.
The statement in (14) reveals that the weights are sandwiched between two logistic functions with midpoints $h(\alpha,k)$ and $h(\alpha,k)+1$ and growth rate $\alpha=-\log(1-\gamma)$. Moreover, when $\alpha k$ is large, $h(\alpha,k)\approx k$, and when $\gamma$ is small, $\alpha\approx\gamma$, which reveals how the RGS parameters directly affect the shape of its weight function. [^8] In fact, if the sampling of the features in RGS is performed with replacement, then all asymptotic results hold with $\alpha$ replaced by $\gamma$. Specifically, as $k$ increases, the function gets translated towards the right; as $\gamma$ increases (corresponding to larger $m$), the function becomes more compressed around its midpoint and thereby more closely approximates the sharp thresholding behavior of FS. Figure 2 illustrates this behavior by plotting $\tilde{w}_{j}^{k,\gamma}$ for three values of $\gamma$: $0.05$, $0.15$, and $0.333$.
Figure 2: Comparison of $\tilde{w}_{j}^{k,\gamma}$ (solid) and its logistic bounds (14) (dashed) from Theorem 4.1 for $\gamma=0.05,0.15,$ and $0.333$, with $k=10$.
Interestingly, the approximation of the weights by a logistic function becomes more exact, differing only in a half-index shift as opposed to a full-index shift, if we also let the selection threshold $k$ tend to infinity. In the following theorem, we introduce a new parameter $d=k-j$, which represents the “offset” from the selection threshold.
@@
The limiting weights $\check{w}_{d}^{\gamma}\coloneqq\lim\limits_{\begin{subarray}{c}k\to\infty\end{subarray}}\lim\limits_{\begin{subarray}{c}m\to\infty,\,p\to\infty,\\ m/p\to\gamma\end{subarray}}w_{k-d}^{k,m,p}$
@@
    have a sigmoidal (e.g., S-shaped) pattern with respect to $d$ [^9] This means that (i) $\check{w}_{d}^{\gamma}\leq\check{w}_{d+1}^{\gamma},\quad d\in\mathbb{Z}$, (ii) $\check{w}_{d}^{\gamma}-\check{w}_{d-1}^{\gamma}\leq\check{w}_{d+1}^{\gamma}-\check{w}_{d}^{\gamma},\quad d<0$, (iii) $\check{w}_{d}^{\gamma}-\check{w}_{d-1}^{\gamma}\geq\check{w}_{d+1}^{\gamma}-\check{w}_{d}^{\gamma},\quad d\geq 0$.
Observe that properties (b) and (d) in Theorem 4.2 both also apply to the logistic function. Finally, since $w_{j}^{k,m,p}\approx\check{w}_{k-j}^{\gamma}$ and $\gamma=1-e^{-\alpha}\approx m/p$ for large $k$, $m$, and $p$, the refined logistic bounds in part (c) of Theorem 4.2 imply the approximation formula we stated earlier as (7).
4.2 Comparisons with Elastic Net

### v3

@@ 3.1 EFS($k,m$) as reweighted OLS
The design matrix satisfies $\mathbf{X}^{\top}\mathbf{X}/n=\mathbf{I}$.
This assumption is unlikely to describe most real datasets, but it allows us to derive exact asymptotic expressions for coefficient behavior and precisely characterize how the randomization from feature subsampling propagates through the FS($k$) algorithm. This enables direct comparisons with the broader regularization landscape, including ridge, lasso (tibshirani1996regression), elastic net (zou2005regularization), LARS (efron2004least), SCAD (fan2001variable), MCP (zhang2010nearly), and SureShrink (donoho1995adapting), all of which have been studied in the orthogonal setting.
Let $\hat{f}_{\operatorname{OLS}}=\sum_{j=1}^{p}\hat{\beta}_{j,p}x_{j,p}$ denote the ordinary least squares (OLS) estimator for $f$. Under Assumption 3.1, its coefficients can be expressed as inner products $\hat{\beta}_{j,p}=\langle\mathbf{y},\mathbf{x}_{\cdot j,p}\rangle$. It turns out that both the FS($k$) and EFS($k,m$) estimators can be written in terms of these OLS coefficients. For ease of notation, we establish an ordering convention for the features $x_{1,p},x_{2,p},\ldots,x_{p,p}$ based on these coefficients.
@@
The inner products $\hat{\beta}_{j,p}=\langle\mathbf{y},\mathbf{x}_{\cdot j,p}\rangle$ are reindexed and sorted by magnitude in descending order,
@@
Note that this ordering is random as it depends on the response noise. For simplicity, we assume that the inequalities in (4) are strict apart from a set of measure zero, which holds whenever the noise variables have a continuous density.
Using Definition 3.1 together with the equivalence of FS($k$) and BSS, we can write the FS($k$) estimator as
@@
which reflects the familiar fact that, in the orthogonal setting, FS($k$) (and BSS) selects the top-$k$ features most correlated with the response. The indicators $\mathbf{1}(j\leq k)$ can be thought of as weights that modulate the OLS solution. Similarly, the $k$-step ensemble estimator has the form
$\hat{f}_{k,m,p}=\sum_{j=1}^{p}w_{j}^{k,m,p}\hat{\beta}_{j,p}x_{j,p},$ (6)
where the weights $w_{1}^{k,m,p},w_{2}^{k,m,p},\ldots,w_{p}^{k,m,p}$ are determined solely by structural parameters ($j,k,m,p$) rather than the actual coefficient values $\hat{\beta}_{j,p}$. As $m,p\rightarrow\infty$ with a fixed scaling ratio, we will show that these weights are closely approximated by a logistic function with midpoint $k+1/2$ and growth rate $\log(1-m/p)$,
@@
3.2 Basic properties
Since $w_{j}^{k,m,p}=\mathbb{P}(x_{j,p}\in\mathcal{M}_{k})$, the weight sequence clearly takes values in $[0,1]$ and sums to $k$, the effective number of variables in the model. Unfortunately, the exact values of the weights involve complicated combinatorics and do not have simple closed-form expressions in general. Nonetheless, one can show that the sequence is non-increasing and, more importantly, derive a recurrence relation in terms of $j$, $k$, and $p$, which serves as our main technical tool for the subsequent results.
Lemma 3.1 (Monotonicity).
For all $j<j^{\prime}$ and $k,m,p\geq 1$, we have $w_{j}^{k,m,p}\geq w_{j^{\prime}}^{k,m,p}$.
@@ 3.2 Basic properties
$w_{j}^{k,m,p}=\frac{\binom{p-j}{m-1}}{\binom{p}{m}}+\frac{\binom{p-j}{m}}{\binom{p}{m}}w_{j}^{k-1,m,p-1}+\left(1-\frac{\binom{p-j+1}{m}}{\binom{p}{m}}\right)w_{j-1}^{k-1,m,p-1},$ (8)
@@
Proof sketch of Lemma 3.1.
Define the events $\mathcal{E}_{1}=\{x_{j,p}\in\mathcal{M}_{k},\,x_{j-1,p}\notin\mathcal{M}_{k}\}$ and $\mathcal{E}_{2}=\{x_{j-1,p}\in\mathcal{M}_{k},\,x_{j,p}\notin\mathcal{M}_{k}\}$. It suffices to show $\mathbb{P}(\mathcal{E}_{1})\leq\mathbb{P}(\mathcal{E}_{2})$.
We construct a measure-preserving map $T$ from $\mathcal{E}_{1}$ to $\mathcal{E}_{2}$ as follows. Given a random seed $\boldsymbol{\mathcal{V}}\in\mathcal{E}_{1}$ where $x_{j,p}$ is selected at step $i$, define $T(\boldsymbol{\mathcal{V}})$ by swapping the roles of $x_{j,p}$ and $x_{j-1,p}$ in the candidate sets from step $i$ onward. Since the algorithm depends only on the ordinal structure of coefficients, this transformation preserves probabilities and maps $\mathcal{E}_{1}$ into $\mathcal{E}_{2}$, establishing the desired inequality. ∎
Proof sketch of Lemma 3.2.
Let $J$ denote the index of the feature selected at the first step. The key observation is that conditioning on $J$ reduces the problem to a smaller instance of EFS.
If $J=j$, then $x_{j,p}\in\mathcal{M}_{k}$ with certainty. If $J\neq j$, the remaining k1k-1 steps are equivalent to running EFS on p1p-1 features with reassigned indices. When $J>j$, the index $j$ remains unchanged, contributing $w_{j}^{k-1,m,p-1}$. When $J<j$, the index shifts to j1j-1, contributing $w_{j-1}^{k-1,m,p-1}$.
Computing $\mathbb{P}(J=j)$, $\mathbb{P}(J>j)$, and $\mathbb{P}(J<j)$ via combinatorics (specifically, the probability that $j$ is the minimum index in a uniformly random $m$-subset) and applying the hockey-stick identity to simplify the resulting sums yields the recurrence. ∎
@@ 3.3 Logistic approximation
The recurrence (8) is three-dimensional (in $j$, $k$, and $p$) and non-homogeneous, making it challenging to find a simple closed-form solution. However, in the scaling limit as $m\to\infty$, $p\to\infty$, and $m/p\to\gamma$ for some constant $\gamma\in[0,1]$, the recurrence converges to a two-dimensional system that further decouples into a one-dimensional subsystem with an explicit analytical solution.
@@
The limiting weights $\tilde{w}_{j}^{k,\gamma}=\lim\limits_{\begin{subarray}{c}m\to\infty,\,p\to\infty,\\ m/p\to\gamma\end{subarray}}w_{j}^{k,m,p}$
@@
    where $h(\alpha,k)=\frac{1}{\alpha}\log(e^{\alpha k}-1)$ and $\alpha\in(0,\infty)$ is defined via $\gamma=1-e^{-\alpha}$.
Proof sketch.
The argument proceeds in four steps.
Step 1. We derive the limiting recurrence. Taking $m,p\to\infty$ with $m/p\to\gamma$ in the coefficients of (8) yields
$\frac{\binom{p-j}{m-1}}{\binom{p}{m}}\to(e^{\alpha}-1)e^{-\alpha j},\quad\frac{\binom{p-j}{m}}{\binom{p}{m}}\to e^{-\alpha j},\quad 1-\frac{\binom{p-j+1}{m}}{\binom{p}{m}}\to 1-e^{-\alpha(j-1)},$
where $\alpha=-\log(1-\gamma)$. This produces the limiting recurrence
$\tilde{w}_{j}^{k,\gamma}=(e^{\alpha}-1)e^{-\alpha j}+e^{-\alpha j}\tilde{w}_{j}^{k-1,\gamma}+(1-e^{-\alpha(j-1)})\tilde{w}_{j-1}^{k-1,\gamma}.$
Existence of the limits follows by induction on $k$, using $\tilde{w}_{j}^{0,\gamma}=0$ as the base case.
Step 2. We decouple the recurrence. Through algebraic manipulation, this two-dimensional recurrence separates into two independent one-dimensional recurrences,
$\tilde{w}_{j}^{k,\gamma}$ $=(e^{-\alpha(j-k)}-e^{-\alpha j})(1-\tilde{w}_{j}^{k-1,\gamma}),$
$\tilde{w}_{j}^{k,\gamma}$ $=1-e^{-\alpha k}-(e^{-\alpha(k-j+1)}-e^{-\alpha k})\tilde{w}_{j-1}^{k,\gamma}.$
Step 3. We solve for the closed form. The first recurrence is a first-order linear relation in $k$ with non-constant coefficients. Solving it directly yields the closed form (9).
Step 4. We establish the bounds. The weights satisfy $\tilde{w}_{j}^{k,\gamma}\leq\tilde{w}_{j}^{k+1,\gamma}$ (more steps can only increase inclusion probability). Combining this monotonicity with the first decoupled recurrence gives the lower bound. The upper bound follows by a symmetric argument using $\tilde{w}_{j}^{k,\gamma}\geq\tilde{w}_{j}^{k-1,\gamma}$. ∎
The bounds in (10) reveal that the weights are sandwiched between two logistic functions with midpoints $h(\alpha,k)$ and $h(\alpha,k)+1$ and common growth rate $\alpha=-\log(1-\gamma)$. When $\alpha k$ is large and $\gamma$ is small, we have $h(\alpha,k)\approx k$ and $\alpha\approx\gamma$, respectively.
These approximations clarify how the EFS($k,m$) parameters affect the shape of the weight function. (If sampling is performed with replacement, all asymptotic results hold with $\alpha$ replaced by $\gamma$.) As $k$ increases, the logistic curve translates rightward. As $\gamma$ increases (larger $m$), the curve becomes steeper, more closely approximating the sharp thresholding of FS($k$). Figure 1 illustrates this behavior for $\gamma\in\{0.05,0.15,0.333\}$.
Figure 1: Comparison of $\tilde{w}_{j}^{k,\gamma}$ (solid) and the logistic approximation (7) (dashed) for $\gamma=0.05,0.15,$ and $0.333$, with $k=10$.
The logistic approximation becomes even more precise, differing only by a half-index shift rather than a full-index shift, when the selection threshold $k$ also tends to infinity. We capture this by introducing the “offset” parameter d=kjd=k-j, measuring distance from the selection threshold.
@@
The limiting weights $\check{w}_{d}^{\gamma}=\lim\limits_{\begin{subarray}{c}k\to\infty\end{subarray}}\lim\limits_{\begin{subarray}{c}m\to\infty,\,p\to\infty,\\ m/p\to\gamma\end{subarray}}w_{k-d}^{k,m,p}$
@@
    have a sigmoidal (S-shaped) pattern with respect to $d$.
Proof sketch.
We establish each part in turn.
Part (a). Using the recurrences from Theorem 3.1, one can show $\tilde{w}_{j+1}^{k,\gamma}-\tilde{w}_{j}^{k-1,\gamma}\geq 0$. For fixed offset d=kjd=k-j, the sequence $\{\tilde{w}_{k-d}^{k,\gamma}\}_{k\geq d+1}$ is therefore bounded and non-decreasing, hence convergent by the monotone convergence theorem. Taking $k\to\infty$ in the closed form (9) with j=kdj=k-d, careful analysis of convergence rates (tail terms decay exponentially) yields (11).
Part (b). Direct calculation using (11) verifies $1-\check{w}_{-(d+1)}^{\gamma}=\check{w}_{d}^{\gamma}$.
Part (c). These follow from taking $k\to\infty$ in the bounds (10).
Part (d). The bounds immediately imply that $\check{w}_{d}^{\gamma}$ is non-decreasing in $d$. Convexity for d<1/2d<-1/2 and concavity for d>1/2d>-1/2 are established by analyzing the second derivative of an extended continuous version of $\check{w}_{d}^{\gamma}$, using a $q$-series identity from coogan2003q. ∎
4 Main results
We now present our main results comparing EFS($k,m$) to FS($k$) in terms of the training error and degrees of freedom decomposition (3).
4.1 Expected training error
Let $\hat{m}$ denote the value of $m$ minimizing the training error $\|\mathbf{y}-\hat{\mathbf{f}}_{k,m,p}\|^{2}$. Since FS($k$) coincides with EFS($k,m$) when $m=p$, the expected training error of EFS($k,\hat{m}$) is trivially no larger than that of FS($k$). What is surprising is that $\hat{m}$ can take values less than $m$ and that we can quantify the gap in expected training error using the asymptotic formulas (10).
Theorem 4.1 (Training error improvement).
Grant Assumption 3.1 and fix $k\geq 1$. Assume $p$ is large enough that
$p\geq C\Bigg(\frac{\sum_{j=1}^{k}\mathbb{E}\big[\hat{\beta}^{2}_{j,p}\big]}{\sum_{j=k+1}^{2k}\mathbb{E}\big[\hat{\beta}^{2}_{j,p}\big]}\Bigg)^{3},$
where $C$ is a positive constant depending only on $k$. Then
$\mathbb{E}\big[\|\mathbf{y}-\hat{\mathbf{f}}_{k,\hat{m},p}\|^{2}\big]\leq\mathbb{E}\big[\|\mathbf{y}-\hat{\mathbf{f}}_{k,p}\|^{2}\big]-\frac{1}{4}\frac{\big(\frac{1}{k}\sum_{j=k+1}^{2k}\mathbb{E}\big[\hat{\beta}^{2}_{j,p}\big]\big)^{2}}{\frac{1}{k}\sum_{j=1}^{k}\mathbb{E}[\hat{\beta}^{2}_{j,p}]}.$ (12)
Proof sketch.
The argument proceeds in four steps.
Step 1. We decompose the training error gap. From (6) and (5), the gap decomposes as
$\|\mathbf{y}-\hat{\mathbf{f}}_{k,p}\|^{2}-\|\mathbf{y}-\hat{\mathbf{f}}_{k,m,p}\|^{2}=\underbrace{\sum_{j=k+1}^{p}\hat{\beta}_{j,p}^{2}\big(1-(1-w_{j}^{k,m,p})^{2}\big)}_{\text{gain from including }j>k}-\underbrace{\sum_{j=1}^{k}\hat{\beta}_{j,p}^{2}(1-w_{j}^{k,m,p})^{2}}_{\text{loss from excluding }j\leq k}.$
Step 2. We apply Chebyshev’s sum inequality. Both $\mathbb{E}[\hat{\beta}_{j,p}^{2}]$ and $w_{j}^{k,m,p}$ are non-increasing in $j$. Chebyshev’s inequality therefore bounds the loss term from above and the gain term from below by products of averages,
$\sum_{j=1}^{k}\mathbb{E}[\hat{\beta}_{j,p}^{2}](1-w_{j}^{k,m,p})^{2}$ $\leq\frac{1}{k}\sum_{j=1}^{k}\mathbb{E}[\hat{\beta}_{j,p}^{2}]\cdot\sum_{j=1}^{k}(1-w_{j}^{k,m,p})^{2},$
$\sum_{j=k+1}^{2k}\mathbb{E}[\hat{\beta}_{j,p}^{2}]\big(1-(1-w_{j}^{k,m,p})^{2}\big)$ $\geq\frac{1}{k}\sum_{j=k+1}^{2k}\mathbb{E}[\hat{\beta}_{j,p}^{2}]\cdot\sum_{j=k+1}^{2k}\big(1-(1-w_{j}^{k,m,p})^{2}\big).$
Step 3. We bound the weight sums using the logistic approximation. The bounds (10) imply $w_{j}^{k,\gamma}\leq 1/2$ for $j\geq k+1$, so $1-(1-w_{j}^{k,\gamma})^{2}\geq\frac{3}{2}w_{j}^{k,\gamma}$. Combined with the lower bound on $w_{j}^{k,\gamma}$, this gives $\sum_{j=k+1}^{2k}(1-(1-w_{j}^{k,\gamma})^{2})\gtrsim 1-\gamma$. A similar analysis yields $\sum_{j=1}^{k}(1-w_{j}^{k,\gamma})^{2}\lesssim(1-\gamma)^{2}$.
Step 4. We optimize over $\gamma$. Let $A=\frac{1}{k}\sum_{j=1}^{k}\mathbb{E}[\hat{\beta}_{j,p}^{2}]$ and $B=\frac{1}{k}\sum_{j=k+1}^{2k}\mathbb{E}[\hat{\beta}_{j,p}^{2}]$. The gap is at least $c_{1}B(1-\gamma)-c_{2}A(1-\gamma)^{2}$ for universal constants $c_{1},c_{2}>0$. Optimizing over $\gamma$ and verifying the condition on $p$ ensures the asymptotic approximations are valid, yields the bound $\frac{1}{4}B^{2}/A$. ∎
The bound (12) is larger when the coefficients $\hat{\beta}_{k+1,p},\hat{\beta}_{k+2,p},\ldots$ decay slowly. This makes intuitive sense, as in such settings spreading the weight budget more evenly across features improves the fit.
4.2 Degrees of freedom
We now show that, under a natural assumption, decreasing the feature subsampling ratio $m/p$ also decreases the degrees of freedom. The key is a decomposition of the degrees of freedom of EFS($k,m$) as a weighted sum over those of FS($j$) for varying $j$.
Theorem 4.2 (Degrees of freedom improvement).
Grant Assumption 3.1. For all $k,m,p\geq 1$, the degrees of freedom satisfy
$\operatorname{df}(\hat{f}_{k,m,p})=\sum_{j=1}^{p}w_{j}^{k,m,p}\left(\operatorname{df}(\hat{f}_{j,p})-\operatorname{df}(\hat{f}_{j-1,p})\right),$ (13)
where $\hat{f}_{0,p}=0$ and $\operatorname{df}(\hat{f}_{0,p})=0$. In particular, if $k\mapsto\operatorname{df}(\hat{f}_{k,p})$ is concave, then $m\mapsto\operatorname{df}(\hat{f}_{k,m,p})$ is increasing.
Proof sketch.
The decomposition (6) implies
$\hat{f}_{k,m,p}=\sum_{j=1}^{p}w_{j}^{k,m,p}(\hat{f}_{j,p}-\hat{f}_{j-1,p}).$
Since degrees of freedom is linear in the estimator, (13) follows immediately.
For the second statement, we establish that the weight sequences form a majorization hierarchy. For $m<m^{\prime}$,
$\sum_{j=1}^{l}w_{j}^{k,m,p}\leq\sum_{j=1}^{l}w_{j}^{k,m^{\prime},p}\quad\text{for all }l=1,\dots,p,$
with equality when $l=p$. The conclusion then follows from a standard majorization inequality (marshall11).
To prove the majorization, we construct a coupling between the EFS processes at levels $m$ and $m^{\prime}$. At each iteration, first draw a subset $B_{m}$ of size $m$, then augment it to $B_{m^{\prime}}\supseteq B_{m}$ of size $m^{\prime}$ by adding $m^{\prime}-m$ additional elements. Let $Z_{k,m,p}(l)=|S_{k,m,p}\cap[l]|$ count selected features with index at most $l$. Note that $\mathbb{E}[Z_{k,m,p}(l)]=\sum_{j=1}^{l}w_{j}^{k,m,p}$.
An inductive argument shows $Z_{k,m^{\prime},p}(l)\geq Z_{k,m,p}(l)$ for all $l$ and $k$. The key insight is that $\min B_{m^{\prime}}\leq\min B_{m}$, so the feature selected with parameter $m^{\prime}$ has index at most that selected with $m$. Careful case analysis completes the induction. ∎
The concavity condition in Theorem 4.2 is likely to hold in low-SNR scenarios where $|\hat{\beta}_{k,p}|\gg|\beta_{k,p}|$. Indeed, in such cases,
$\frac{\sigma^{2}}{n}\left(\operatorname{df}(\hat{f}_{k,p})-\operatorname{df}(\hat{f}_{k-1,p})\right)=\frac{\sigma^{2}}{n}\operatorname{df}\big(\hat{\beta}_{k,p}x_{k,p}\big)=\mathbb{E}\big[\hat{\beta}_{k,p}\big(\hat{\beta}_{k,p}-\beta_{k,p}\big)\big]\approx\mathbb{E}\big[\hat{\beta}_{k,p}^{2}\big],$ (14)
which is non-increasing in $k$ by Definition 3.1. Empirical support for the concavity of FS($k$) degrees of freedom, even in non-orthogonal settings, appears in Figures 1 and 4 of hastie2020extended and Figure 7 of mentch2020randomization.
Remark.
For a constant model ($\boldsymbol{\beta}=\mathbf{0}$), the approximation in (14) becomes exact.
When concavity fails, the degrees of freedom can be non-monotonic in $m$. This happens, for instance, when applying FS($k$) and EFS($k,m$) to data generated from an exactly $k$-sparse model.
Proposition 4.1 (Non-monotonic degrees of freedom).
Grant Assumption 3.1 and suppose $\boldsymbol{\beta}$ is exactly $k$-sparse with all nonzero coefficients equal to 1, and that the noise is Gaussian. Then
$\lim\limits_{\begin{subarray}{c}m,\,p,\,n\to\infty,\\ m/p\to\gamma\end{subarray}}\frac{\operatorname{df}(\hat{f}_{k,m,p})-\operatorname{df}(\hat{f}_{k,p})}{\log p}\geq C,$
where $C>0$ depends only on $\gamma$ and $k$.
Proof sketch.
Because of (13), we know that $\operatorname{df}(\hat{f}_{k,m,p})\geq w_{j}^{k,m,p}\operatorname{df}(\hat{\beta}_{k+1,p}x_{k+1,p})$. It hence suffices to show that $\operatorname{df}(\hat{\beta}_{k+1,p}x_{k+1,p})\approx\mathbb{E}[\hat{\beta}_{k+1,p}^{2}]$ is of order $\log p$. This follows from the fact that $\hat{\beta}_{k+1,p}x_{k+1,p}$ is essentially the maximum of pkp-k independent $\chi_{1}^{2}$ random variables. ∎
5 Further comparisons
5.1 Comparison with elastic net